OpenStudy (anonymous):

Calculate the force that earth's gravity exerts on a 1 kg mass at the following locations: a) On the surface of the earth b) 3 km above earth's surface c) 300 km above earths surface d) 30,000 km above earths surface Please help me with the answers, I wasnt here when we learned this.

OpenStudy (anonymous):

we all know when the exerts force on the body it becomes weight so and weight is a force so force = m*g whereas g= G * M / d^2 G= gravitational constent M = mass of the earth d = distence of the earth from centre then the above formula becomes f = G*M*m/d^2 as f= mg m*g= G*M*m/d^2 m on both sides will cancel each other g = G*M/d^2 1 where value of G =6.673 * 10^-11 Nm^2 value of M = 5.98 * 10 ^24 kg and value of d at surface of earth is = 6.38 * 10^ 6 you can find the value by using this formula g=g = G*M/d^2

OpenStudy (shamim):

3km above earth surface g=GM/(R+h)^2 h=3000m

OpenStudy (anonymous):

R = 6400 km

OpenStudy (anonymous):

Ohhhh I'll use R=6400 with shamims equation?